# Simplicity Prevails: The Emergence of Generalizable AIGI Detection in Visual Foundation Models

### v1

@@
Abstract
While specialized detectors for AI-Generated Images (AIGI) achieve near-perfect accuracy on curated benchmarks, they suffer from a dramatic performance collapse in realistic, in-the-wild scenarios. In this work, we demonstrate that simplicity prevails over complex architectural designs. A simple linear classifier trained on the frozen features of modern Vision Foundation Models , including Perception Encoder, MetaCLIP 2, and DINOv3, establishes a new state-of-the-art. Through a comprehensive evaluation spanning traditional benchmarks, unseen generators, and challenging in-the-wild distributions, we show that this baseline not only matches specialized detectors on standard benchmarks but also decisively outperforms them on in-the-wild datasets, boosting accuracy by striking margins of over 30%. We posit that this superior capability is an emergent property driven by the massive scale of pre-training data containing synthetic content. We trace the source of this capability to two distinct manifestations of data exposure: Vision-Language Models internalize an explicit semantic concept of forgery, while Self-Supervised Learning models implicitly acquire discriminative forensic features from the pretraining data. However, we also reveal persistent limitations: these models suffer from performance degradation under recapture and transmission, remain blind to VAE reconstruction and localized editing. We conclude by advocating for a paradigm shift in AI forensics, moving from overfitting on static benchmarks to harnessing the evolving world knowledge of foundation models for real-world reliability.
Machine Learning, ICML
1 Introduction
@@
The rapid evolution of generative models, such as Midjourney, Stable Diffusion (Rombach et al., 2021) and Nano Banana(Team et al., 2023), has ushered in a new era of content creation, synthesizing photorealistic images that challenge the boundaries of visual authenticity. While empowering creativity, this technological leap simultaneously introduces profound threats to information integrity, fueling the proliferation of misinformation and deepfakes. In response, the forensics community has largely favored a specialized approach: crafting detectors with increasingly complex module tailored to specific artifacts, such as frequency anomalies or noise residuals (Wang et al., 2020; Ju et al., 2022). While these specialized detectors achieve near-perfect accuracy on curated benchmarks, they often suffer from a dramatic performance collapse in realistic scenarios. Recent studies, such as the Chameleon benchmark (Yan et al., 2024a), reveal that detectors excelling in controlled environments frequently degrade to 60%–70% accuracy when deployed ‘in-the-wild’. This fragility suggests that relying on hand-crafted inductive biases may be a dead end in the face of rapidly evolving generative distributions.
@@
Echoing Sutton’s “Bitter Lesson”, we present a finding that challenges this trend: simplicity prevails. We demonstrate that a simple linear classifier, trained on top of frozen features from modern Vision Foundation Models (VFMs), achieves state-of-the-art performance. We define modern VFMs as the latest generation of encoders, such as Perception Encoder (PE) (Bolya et al., 2025), MetaCLIP 2 (Chuang et al., 2025), and DINOv3 (Siméoni et al., 2025), which have been exposed to billion-scale web data. Our evaluation spans three distinct protocols: standard benchmarks (e.g., GenImage (Zhu et al., 2023)), datasets from the latest unseen generators (e.g., AIGIHolmes (Zhou et al., 2025b), AIGI-Now (Chen et al., 2025a)), and challenging in-the-wild distributions (e.g., Chameleon(Yan et al., 2024a), WildRF(Cavia et al., 2024)). Across all settings, this simple baseline matches or decisively outperforms specialized detectors, boosting accuracy by margins of over 30% in the most difficult in-the-wild scenarios.
@@
We posit that this superior capability is not an architectural innovation for forensics but an emergent property derived from the massive scale of pre-training data. Through rigorous analysis of the Common Crawl index [^1] Common Crawl Index Server https://index.commoncrawl.org/, we observe an exponential explosion of generative content in web corpora starting from 2022. Modern VFMs, trained on this evolving data landscape, have inadvertently internalized the necessary concepts to distinguish synthetic content. We trace the source of this capability to two distinct manifestations of data exposure. For Vision-Language Models, the co-occurrence of synthetic images and textual descriptions leads to explicit concept injection, where the model aligns synthetic visuals with high-level forgery concepts. For Self-Supervised Learning (SSL) models like DINOv3, the capability emerges from implicit distribution fitting, where the model encodes the distinct low-level signatures of the generative manifold in pretraining data.
@@
Despite these significant gains, our analysis delineates the boundaries of this paradigm. We find that modern VFMs remain blind to pure reconstruction artifacts and struggle with localized editing, while also suffering from performance degradation under aggressive real-world transmission and screen recapture. This suggests that while data scale has effectively solved the generalizability problem for fully synthetic content, the localization of fine-grained manipulation remains an open challenge. Therefore, we advocate for a paradigm shift in AI forensics: moving away from designing specialized architectures to chase benchmark metrics—since simple backbones already suffice for global detection—and instead focusing on how to harness and refine the evolving world knowledge of foundation models to address these fine-grained forensic tasks.
@@
  • We establish the superiority of simple linear probes on modern VFMs over specialized detectors. Across a comprehensive evaluation spanning standard benchmarks, previously unseen generators, and challenging in-the-wild distributions, our baseline achieves SOTA performance, outperforming specialized methods by margins of over 30% in realistic scenarios.
  • We trace the source of this capability to pre-training data exposure rather than architectural design. We identify two emergent mechanisms: explicit concept injection in Vision Language Models and implicit distribution fitting in Self-Supervised Learning Models.
  • We delineate the boundaries of this paradigm, revealing persistent blind spots in pure reconstruction and localized editing. We conclude that future research should shift focus from designing complex architectures to harnessing the evolving world knowledge of foundation models for fine-grained forensic tasks.
Table 1: Performance on GenImage Benchmark. All detectors are trained on Stable Diffusion v1.4 and evaluated on unseen generators. Accuracy is averaged over real and fake classes. Best results in bold.
@@ 2 Related Works
Data-Driven Specialized Detectors. With the dominance of Diffusion Models, the focus shifted from detecting CNN-specific upsampling artifacts (e.g., checkerboard patterns in GANs (Wang et al., 2020)) to identifying the unique traces of the diffusion process. Researchers proposed reconstructing input images to isolate generative errors: DIRE (Wang et al., 2023) leverages the reconstruction residual from a pre-trained diffusion model as a forensic signal, while DRCT (Chen et al., 2024a) refines this by analyzing the discrepancies between real-real and fake-fake reconstruction pairs. Others focused on improving generalization: SAFE (Li et al., 2025b) introduces artifact-preserving augmentations to decouple semantic content from forensic traces. notably, DDA (Chen et al., 2025b) targets the shared VAE decoder inherent to Latent Diffusion Models, explicitly aligning the detector with VAE reconstruction patterns to achieve broader generalization across different LDM-based generators.
@@
The Foundation Model Era. The introduction of UnivFD  (Ojha et al., 2023) marked a pivotal turning point. Ojha et al. revealed that training a linear layer on top of the frozen feature space of a pre-trained Vision-Language Model (specifically CLIP (Radford et al., 2021)) yields significantly better generalization than training CNNs from scratch. This discovery spurred a wave of research leveraging Vision Foundation Models (VFMs) as backbones. Subsequent works, such as Effort(Yan et al., 2024b), AIDE (Yan et al., 2024a), OMAT(Zhou et al., 2025a) and DDA(Chen et al., 2025b), have explored various strategies to adapt CLIP or DINOv2 (Oquab et al., 2023) for forensics, including prompt tuning, adapter modules, fusing frequency-domain information and training hard samples. These methods currently represent the state-of-the-art, yet as our experiments show, they still struggle to maintain robustness in unconstrained, in-the-wild scenarios compared to the raw capabilities of the latest VFMs. (See Appendix A for a comprehensive survey of the legacy backbones still prevalent in these recent methods.)
3 Simplicity Prevails: Benchmarking Modern VFMs
@@
To empirically validate our hypothesis that the generalization capability of AIGI detection stems from the scale of pre-training data rather than complex architectural designs, we conduct a comprehensive comparative analysis. We pit simple linear classifiers trained on modern Vision Foundation Models (VFMs) against a wide array of state-of-the-art specialized detectors. Our evaluation protocol is designed to be rigorous and progressively challenging, spanning three distinct scenarios: standard academic benchmarks, realistic in-the-wild distributions, and unseen next-generation generative models. In this section, we detail our experimental setup and present the performance comparisons that establish the superiority of modern VFM baselines.
@@
Table 2: Performance on In-the-Wild Benchmarks. Evaluation on Chameleon, WildRF, SocialRF, and CommunityAI datasets. Accuracy is averaged over real and fake classes. Best results in bold.
@@
3.1 Experimental Setup
Evaluation Benchmarks. To rigorously assess generalization, we organize our evaluation into three progressively challenging categories. (1) Standard Benchmarks: We use GenImage (Zhu et al., 2023), a widely adopted benchmark comprising images from 8 generators (e.g., Stable Diffusion, Midjourney). Following standard protocols (Ojha et al., 2023), we use the Stable Diffusion v1.4 subset for training and the remaining subsets for testing. (2) In-the-Wild Datasets: We evaluate on Chameleon (Yan et al., 2024a), WildRF (Cavia et al., 2024), SocialRF and CommunityAI (Li et al., 2025d). These datasets are collected from social media and internet forums, featuring diverse, unconstrained post-processing and unknown generative sources, representing a realistic detection scenario. (3) Unseen Generators: We employ AIGIHolmes (Zhou et al., 2025b) and AIGI-Now (Chen et al., 2025a), recent benchmarks containing images from state-of-the-art generators released after 2024, including close source generators like Nano Banana, GPT4o and FLUX-Pro. These serve as a strict test for generalization to unseen distributions.
Specialized Detectors. We compare against a comprehensive suite of state-of-the-art forensic methods, spanning three categories: (1) Specialized Detectors: Artifact-based methods like CNNSpot (Wang et al., 2020), FreqNet (Tan et al., 2024a), and NPR (Tan et al., 2024b), as well as recent VFM-based adapters such as UnivFD (Ojha et al., 2023), OMAT (Zhou et al., 2025a), Effort (Yan et al., 2024b) and Dual-Data-Alignment (DDA)(Chen et al., 2025b). (2) Former VFM Baselines: To explicitly evaluate the impact of pre-training data evolution, we include earlier generations of foundation models, such as the original OpenAI CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), Meta CLIP (Xu et al., 2023a) and DINOv2 (Oquab et al., 2023). For DDA, we utilize the official pre-trained weights provided by the authors, as its core contribution involves a specialized training pipeline with VAE-reconstructed data alignment. All other methods are trained on GenImage stable_diffusion_v_1_4 training set for fair comparison.
Modern VFM Baselines. To test our “Simplicity Prevails” hypothesis, we select a representative set of modern Vision Foundation Models as frozen feature extractors. These include Vision-Language Models (CLIP (Radford et al., 2021), SigLIP2 (Tschannen et al., 2025), MetaCLIP 2 (Chuang et al., 2025), Perception Encoder (Bolya et al., 2025)) and Self-Supervised Models (DINOv2 (Oquab et al., 2023), DINOv3 (Siméoni et al., 2025)). We attach a simple linear layer to the pooled output features of these backbones. Detailed specifications of model architectures and pre-training datasets are provided in Appendix C.
Implementation Details. All our VFM baselines are trained solely on the GenImage (SD v1.4) training set. We keep the backbone completely frozen and only update the linear head. We use the AdamW optimizer with a learning rate of $1e^{-3}$ and a batch size of 128 for 2 epoch. Images are resized and center-cropped to the native resolution of each model without any additional data augmentation.
3.2 Performance on Standard Benchmarks
@@
Table 3: Generalization on AIGIHolmes. Evaluation on advanced auto-regressive and diffusion-transformer generative models. Accuracy is averaged over real and fake classes. Best results in bold.

### v2

@@
Abstract.
Specialized detectors for AI-generated images (AIGI) often achieve near-perfect accuracy on curated benchmarks, yet their performance degrades substantially in realistic, in-the-wild scenarios. In this work, we show that frozen features from modern Vision Foundation Models (VFMs), combined with a lightweight classifier, form a remarkably strong baseline for generalizable AIGI detection. Using representative modern encoders, including Perception Encoder, MetaCLIP 2, and DINOv3, we conduct a comprehensive evaluation across standard benchmarks, recent unseen generators, and challenging in-the-wild distributions. Across these settings, this simple baseline consistently matches or outperforms recent specialized detectors, with particularly large gains in realistic scenarios.
We further investigate why this simple setup is so effective. Our analyses provide converging evidence that the strong forensic separability of modern VFMs is closely related to their exposure to synthetic web content during pre-training. In Vision-Language Models, this manifests as semantic alignment with forgery-related concepts, while in Self-Supervised Learning models it appears as implicit discrimination of generative distributions. Although a fully controlled pre-training study is beyond the scope of this work, multiple complementary analyses support this interpretation. We also identify important limitations. While modern VFMs are highly effective for global AIGI detection, they remain vulnerable to severe transmission degradation and perform poorly on pure VAE reconstruction and localized editing. Overall, our results suggest that progress in generalizable AIGI detection may depend more on preserving and leveraging strong pretrained representations than on increasingly complex task-specific forensic designs.
AI-generated image detection, multimedia forensics, vision foundation models
copyright: noneconference: Proceedings of the 34th ACM International Conference on Multimedia; October 2026; Amsterdam, The Netherlandsbooktitle: Proceedings of the 34th ACM International Conference on Multimedia (ACM $M$ ’26), October 2026, Amsterdam, The Netherlandsjournalyear: 2026doi: TBDisbn: TBDccs: Computing methodologies Computer visionccs: Computing methodologies Artificial intelligenceccs: Information systems Multimedia content analysis
1. Introduction
@@
The rapid evolution of generative models, such as Midjourney, Stable Diffusion (Rombach et al., 2021) and Nano Banana(Team et al., 2023), has ushered in a new era of content creation, synthesizing photorealistic images that challenge the boundaries of visual authenticity. While empowering creativity, this technological leap simultaneously introduces profound threats to information integrity, fueling the proliferation of misinformation. In response, the forensics community has largely favored a specialized approach: crafting detectors with increasingly complex module tailored to specific artifacts, such as frequency anomalies or noise residuals (Wang et al., 2020; Ju et al., 2022). While these specialized detectors achieve near-perfect accuracy on curated benchmarks, they often suffer from a dramatic performance collapse in realistic scenarios. Recent studies, such as the Chameleon benchmark (Yan et al., 2024a), reveal that detectors excelling in controlled environments frequently degrade to 60%–70% accuracy when deployed ‘in-the-wild’. This fragility suggests that relying on hand-crafted inductive biases may be a dead end in the face of rapidly evolving generative distributions.
Figure 1. The Surge of Generative Data in Web Corpora. We track the number of indexed URLs from major open-source AI generation platforms (Civitai and Liblib) within Common Crawl snapshots from 2022 to 2025.
@@
Echoing Sutton’s “Bitter Lesson”, we revisit this trend from a different perspective: rather than asking how to design increasingly specialized forensic modules, we ask how far one can go by directly leveraging the frozen representations of modern Vision Foundation Models (VFMs). We show that a simple linear classifier, trained on top of frozen features from recent encoders such as Perception Encoder (PE) (Bolya et al., 2025), MetaCLIP 2 (Chuang et al., 2025), and DINOv3 (Siméoni et al., 2025), provides a remarkably strong baseline for generalizable AIGI detection. We define modern VFMs as the latest generation of encoders trained on large-scale and evolving web corpora. Our evaluation spans three distinct protocols: standard benchmarks (e.g., GenImage (Zhu et al., 2023)), datasets from the latest unseen generators (e.g., AIGIHolmes (Zhou et al., 2025b), AIGI-Now (Chen et al., 2025a)), and challenging in-the-wild distributions (e.g., Chameleon (Yan et al., 2024a), WildRF (Cavia et al., 2024)). Across all settings, this simple baseline consistently matches or outperforms recent specialized detectors, with the largest gains appearing in the most challenging in-the-wild scenarios.
@@
We further investigate why such a simple setup is so effective. Rather than attributing the gains to forensic-specific architectural innovation, we argue that they are closely related to an emergent property of large-scale pre-training on evolving web data. As visualized in Figure 1, our analysis of the Common Crawl index reveals a rapid increase in generative content from major communities such as Civitai and Liblib starting in 2023. This trend suggests that modern VFMs are increasingly likely to encounter synthetic content during pre-training, and may therefore internalize useful cues for distinguishing generated images from real ones. We characterize this capability through two distinct manifestations of data exposure. For Vision-Language Models, the co-occurrence of synthetic images and textual descriptions can lead to explicit concept injection, where synthetic visuals become aligned with high-level forgery-related concepts. For Self-Supervised Learning (SSL) models such as DINOv3, the capability appears instead as implicit distribution fitting, where the model captures low-level regularities associated with the generative manifold through pre-training data exposure.
@@
At the same time, our analysis also clarifies the boundaries of this paradigm. We find that modern VFMs remain blind to pure reconstruction artifacts and struggle with localized editing, while also suffering from noticeable degradation under aggressive real-world transmission and screen recapture. These findings suggest that large-scale pre-training has substantially improved the generalizability of global detection for fully synthetic content, but that the localization of fine-grained manipulation remains an open challenge. Complementing these observations, our experiments on backbone replacement and lightweight LoRA fine-tuning suggest that the dominant factor behind performance is still the strength of the pretrained representation itself, while adaptation mainly serves to better exploit this foundation. Therefore, rather than viewing increasingly specialized detector design as the default path forward, we argue that future progress in AI forensics may depend more on how to preserve, harness, and refine the evolving representations of foundation models for fine-grained forensic reasoning.
@@
  • We show that frozen features from modern Vision Foundation Models, combined with a lightweight classifier, constitute a remarkably strong baseline for generalizable AIGI detection. Across standard benchmarks, recent unseen generators, and challenging in-the-wild distributions, this simple setup consistently matches or outperforms recent specialized detectors, with especially large gains in realistic scenarios.
  • We provide converging evidence that this capability is closely related to pre-training data exposure rather than forensic-specific architectural design. In particular, we identify two complementary manifestations: semantic alignment with forgery-related concepts in Vision-Language Models, and implicit discrimination of generative distributions in Self-Supervised Learning models.
  • We validate the “Bitter Lesson” in AI forensics and delineate the boundaries of this paradigm. Through rigorous ablation, we demonstrate that attaching complex forensic heads or applying fine-tuning (e.g., LoRA) actively degrades the generic representations of VFMs. While these frozen generic features solve global detection, they remain blind to pure VAE reconstruction and localized editing, urging future research to harness rather than over-engineer foundation models.
Table 1. Performance on GenImage Benchmark. All detectors are trained on Stable Diffusion v1.4 and evaluated on unseen generators. Accuracy is averaged over real and fake classes. Best results in bold.
@@ 2. Related Works
Data-Driven Specialized Detectors. With the dominance of Diffusion Models, the focus shifted from detecting CNN-specific upsampling artifacts (e.g., checkerboard patterns in GANs (Wang et al., 2020)) to identifying the unique traces of the diffusion process. Researchers proposed reconstructing input images to isolate generative errors: DIRE (Wang et al., 2023) leverages the reconstruction residual from a pre-trained diffusion model as a forensic signal, while DRCT (Chen et al., 2024) refines this by analyzing the discrepancies between real-real and fake-fake reconstruction pairs. Others focused on improving generalization: SAFE (Li et al., 2025b) introduces artifact-preserving augmentations to decouple semantic content from forensic traces. notably, DDA (Chen et al., 2025b) targets the shared VAE decoder inherent to Latent Diffusion Models, explicitly aligning the detector with VAE reconstruction patterns to achieve broader generalization across different LDM-based generators.
@@
The Foundation Model Era. The introduction of UnivFD  (Ojha et al., 2023) marked a pivotal turning point. Ojha et al. revealed that training a linear layer on top of the frozen feature space of a pre-trained Vision-Language Model (specifically CLIP (Radford et al., 2021)) yields significantly better generalization than training CNNs from scratch. This discovery spurred a wave of research leveraging Vision Foundation Models (VFMs) as backbones. Subsequent works, such as Effort(Yan et al., 2024b), AIDE (Yan et al., 2024a), OMAT(Zhou et al., 2025a) and DDA(Chen et al., 2025b), have explored various strategies to adapt CLIP or DINOv2 (Oquab et al., 2023) for forensics, including prompt tuning, adapter modules, fusing frequency-domain information and training hard samples. These methods currently represent the state-of-the-art, yet as our experiments show, they still struggle to maintain robustness in unconstrained, in-the-wild scenarios compared to the raw capabilities of the latest VFMs. (See Appendix for a comprehensive survey of the legacy backbones still prevalent in these recent methods.)
3. Simplicity Prevails: Benchmarking Modern VFMs
@@
To empirically validate our hypothesis that the generalization capability of AIGI detection stems from the scale of pre-training data rather than complex architectural designs, we conduct a comprehensive comparative analysis. We pit simple linear classifiers trained on modern Vision Foundation Models (VFMs) against a wide array of state-of-the-art specialized detectors. Our evaluation protocol is designed to be rigorous and progressively challenging, spanning three distinct scenarios: standard academic benchmarks, realistic in-the-wild distributions, and unseen next-generation generative models.
3.1. Experimental Setup
Evaluation Benchmarks. To rigorously assess generalization, we organize our evaluation into three progressively challenging categories. (1) Standard Benchmarks: We use GenImage (Zhu et al., 2023), a widely adopted benchmark comprising images from 8 generators (e.g., Stable Diffusion, Midjourney). Following standard protocols (Ojha et al., 2023), we use the Stable Diffusion v1.4 subset for training and the remaining subsets for testing. (2) In-the-Wild Datasets: We evaluate on Chameleon (Yan et al., 2024a), WildRF (Cavia et al., 2024), SocialRF and CommunityAI (Li et al., 2025c). These datasets are collected from social media and internet forums, featuring diverse, unconstrained post-processing and unknown generative sources, representing a realistic detection scenario. (3) Unseen Generators: We employ AIGIHolmes (Zhou et al., 2025b) and AIGI-Now (Chen et al., 2025a), recent benchmarks containing images from state-of-the-art generators released after 2024, including closed-source generators like Nano Banana, GPT4o and FLUX-Pro. These serve as a strict test for generalization to unseen distributions.
Specialized Detectors. We compare against a comprehensive suite of state-of-the-art forensic methods, spanning three categories: (1) Specialized Detectors: Artifact-based methods like CNNSpot (Wang et al., 2020), FreqNet (Tan et al., 2024a), and NPR (Tan et al., 2024b), as well as recent VFM-based adapters such as UnivFD (Ojha et al., 2023), OMAT (Zhou et al., 2025a), Effort (Yan et al., 2024b) and Dual-Data-Alignment (DDA)(Chen et al., 2025b). (2) Former VFM Baselines: To explicitly evaluate the impact of pre-training data evolution, we include earlier generations of foundation models, such as the original OpenAI CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), Meta CLIP (Xu et al., 2023) and DINOv2 (Oquab et al., 2023). For DDA, we utilize the official pre-trained weights provided by the authors, as its core contribution involves a specialized training pipeline with VAE-reconstructed data alignment. All other methods are trained on GenImage SDv1.4 training set for fair comparison.
Modern VFM Baselines. To test our “Simplicity Prevails” hypothesis, we select a representative set of modern Vision Foundation Models as frozen feature extractors. These include Vision-Language Models (SigLIP2 (Tschannen et al., 2025), MetaCLIP 2 (Chuang et al., 2025), Perception Encoder (Bolya et al., 2025)) and Self-Supervised Models (DINOv3 (Siméoni et al., 2025)). We attach a simple linear layer to the pooled output features of these backbones. Detailed specifications of model architectures and pre-training datasets are provided in Appendix.
@@ 3.1. Experimental Setup
Table 2. Performance on In-the-Wild Benchmarks. Evaluation on Chameleon, WildRF, SocialRF, and CommunityAI datasets. Accuracy is averaged over real and fake classes. Best results in bold.
@@
Implementation Details. All our VFM baselines are trained solely on the GenImage (SD v1.4) training set. We keep the backbone completely frozen and only update the linear head. We use the AdamW optimizer with a learning rate of $1e^{-3}$ and a batch size of 128 for 2 epoch. Images are resized and center-cropped to the native resolution of each model without any additional data augmentation. Notably, PE uses a ViT-L/14 backbone at 336px resolution, which is comparable in scale to several strong VFM-based baselines and recent detectors built on CLIP-L style encoders.
3.2. Performance on Standard Benchmarks
@@
3.3. The Collapse of SOTA in the Wild
While standard benchmarks provide a controlled environment, real-world deployment involves diverse, unconstrained data distributions. To evaluate this, we test on four challenging in-the-wild datasets: Chameleon, WildRF, Social-RF, and CommunityAI. The results, summarized in Table 2, reveal a stark contrast. Most specialized detectors suffer a catastrophic performance collapse. Methods like NPR, LaDeDa, and even recent techniques like OMAT degrade to near-random performance, primarily due to a failure to recognize diverse fake samples. The only exception among specialized methods is DDA, which maintains a respectable average accuracy of 85.0%, due to its alignment with the robust VAE decoder patterns shared across latent diffusion models by training on DINOv2 with VAE reconstruct data.
Table 3. Generalization on AIGIHolmes. Evaluation on advanced auto-regressive and diffusion-transformer generative models. Accuracy is averaged over real and fake classes. Best results in bold.